4. Find the convolution of the following functions a. f(t)=t g(t) = sin 3t b. f(t)=é g(t)=cos2t
Answer #1

We have convolution theorem, t (f×g) t = f flus gCt-u) du la) f(t) = t and glt) and glt) = sin(3t) su ܚܘ ܝ u. Sin C3t-3u) du(b) f(t)=et and g(t) = cos(at) t flt) *g(t) = (f*8)t = Je. eu.cos(at-su) da t t =[e T. cos Cat-w) e. 2 sin (2t-24) du - t tTherefore, *-}. [ ** (v3), emington t -2 Sin (at-24) + cos (2t - 2u o 5 et - 2 sin (at) cos (26) = 5 t a sin (at) cos(at) = e